### v2

# A transference principle for involution-invariant functional Hilbert spaces

Abstract.
Let be an affine-linear involution such that and let be two domains in Let be a -invariant -proper map such that is affine-linear and let be a -invariant reproducing kernel Hilbert space of complex-valued holomorphic functions on It is shown that the space endowed with the norm is a reproducing kernel Hilbert space and the linear mapping defined by is a unitary from onto Moreover, a neat formula for the reproducing kernel of in terms of the reproducing kernel of is given. The above scheme is applicable to symmetrized bidisc, tetrablock, -dimensional fat Hartogs triangle and -dimensional egg domain. Although some of these are known, this allows us to obtain an analog of von Neumannβs inequality for contractive tuples naturally associated with these domains.
Key words and phrases:
proper map, involution, symmetrized bidisc, tetrablock, fat Hartogs triangle, egg domain, Hardy space2020 Mathematics Subject Classification:
Primary 32A36, 47A13; Secondary 32A10, 46E221. Introduction and statement of the main theorem
In this paper, we address the problem of constructing a Hardy space of a given domain in the -dimesional Euclidean space (see [10, 12, 20, 21]). In particular, we provide an alternate approach to this problem based on the theory of reproducing kernel Hilbert spaces. Apart from a motivation from the function-theory, the existence of a reasonable Hardy space (cf. conditions (a) and (b) on [12, p.Β 2172]) allows us to obtain an analog of von Neumannβs inequality for a class of operator tuples naturally associated with the domain in question (cf. [9, TheoremΒ 2.1]). In simple words, the main result of this paper allows us to construct a Hardy space from a known Hardy space of a domain in various situations in which is a -proper image of Our construction is based on a circle of ideas in [13, 20, 26], and it is so general that it also applies to Bergman spaces (see [19]). Another important aspect of this principle is that it gives a transformation formula relating the reproducing kernels in question.
For a set denotes the cardinality of For a positive integer let denote the -fold cartesian product of the complex plane The unit polydisc centred at the origin in is denoted by A map is affine-linear if there exists a linear operator on and such that Let be a domain in that is, a nonempty open connected subset of . The set of zeros of a function is denoted by Let be a domain in The Jacobian of a holomorphic map is the determinant of the Jacobian matrix of and it is denoted by A continuous map is proper if pre-image under of any compact subset of is compact. Let be a proper holomorphic map. It is noticed in [24, SectionΒ 1.2] that there exists a positive integer (to be referred to as the multiplicity of ) such that
For short, we refer to a proper map of multiplicity as -proper map. A continuous map is said to be an involution if We reserve the notation for the set of fixed points of We say that is -invariant if for every The map is said to be -invariant if is -invariant and
Remark 1.1.
Let be domains in and be a holomorphic -proper map. It was shown in the proof of [15, PropositionΒ 2.3] that there exists a holomorphic involution such that
| is not the identity map, and | (2) |
To see the uniqueness, let be two holomorphic involutions satisfying (2). Let and note that If then and hence by (1), This contradiction shows that However, being -proper, has preimages under Consequently,
Let be a domain in The linear space of complex-valued holomorphic functions on is denoted by If is bounded, then we reserve the notation for the Banach algebra of complex-valued bounded holomorphic functions on endowed with the supremum norm By a functional Hilbert space , we understand a reproducing kernel Hilbert space (for short, RKHS) of complex-valued holomorphic functions on (for the definition of RKHS, see [22, DefinitionΒ 1.1]). For any functional Hilbert space there exists a positive semi-definite kernel such that
(refer to [22] for the basic theory of RKHS). A multiplier of is a function such that for every The multiplier algebra is the set of multipliers of By the closed graph theorem and (3), any multiplier yields the operator of multiplication by where denotes the -algebra of bounded linear operators on a Hilbert space Moreover, the operator norm makes a unital subalgebra of that is closed in the weak operator topology (see [22, Corollary 5.24]).
We now define a class of functional Hilbert spaces central to this paper (cf. [7, p.Β 753]). Let be a -invariant domain. A functional Hilbert space is said to be -invariant if for every and
Define the subspaces (cf. [13, p.Β 2]) and of by
Remark 1.2.
Assume that is a -invariant functional Hilbert space. Then and are closed subspaces of This follows from the fact that the convergence of a sequence in an RKHS implies its pointwise convergence. Moreover, and the reproducing kernels and of and respectively, are given by
A general version of the decomposition above has been recorded in [7, p.Β 754], while the βmoreoverβ part may be easily derived from (3) and the fact that any RKHS admits a unique reproducing kernel.
The following is the main result of this paper (a number of similar instances of this have been discussed in the literature; [7, 13, 14, 20, 26]).
Theorem 1.3.
Let be an affine-linear involution and let be two domains in . Let be a -invariant -proper map and let be a -invariant functional Hilbert space with the reproducing kernel Then, the linear space
endowed with the norm
is a Hilbert space. Assume further that and is affine-linear. Then, for the linear mapping defined by we have the following statements
is a unitary from onto
is a reproducing kernel Hilbert space with the reproducing kernel given by
Remark 1.4.
We make several remarks for future reference.
Since is proper, (see the first paragraph on [23, p.Β 301]).
Since is -invariant and by the chain rule, and hence is nonconstant.
Since the zero set of has empty interior (by (a)), by the open mapping theorem, defines a norm on the linear space
Since is -invariant, by (b), maps into .
It may happen that Indeed, if is the Hardy space of and then
The formula (10) can be seen as a variant of Bellβs formula for -proper maps (cf. [4, TheoremΒ 1], [26, CorollaryΒ 1]). One of the ingredients of the proof of TheoremΒ 1.3 is a variant of the removable singularity theorem for functional Hilbert spaces (see LemmaΒ 2.1). Another ingredient in the proof of TheoremΒ 1.3 is an abstract version of the Waring-Lagrange theorem (see [2, TheoremΒ 7.1]) for -invariant holomorphic functions see TheoremΒ 2.7 For another variant of the Waring-Lagrange theorem, see [6, Theorem 3.1 & SubsectionΒ 3.1.1]. It is worth noting that a rather special case of TheoremΒ 2.7 (when is linear) can be deduced from [6, Theorem 3.1 & SubsectionΒ 3.1.1] and [13, PropositionΒ 2.2]. On the other hand, our proof of TheoremΒ 2.7 is of analytic flavor and it capitalizes on one of variants of the classical Riemann removable singularity theorem (for short, RRST). Note that the idea of applying -version of RRST (see [18, TheoremΒ 4.2.9]) in obtaining a transformation rule for Bergman kernels appears in [4, 26]. This argument falls short in our case as the spaces under consideration are abstract functional Hilbert spaces. In the first half of SectionΒ 3, we apply TheoremΒ 1.3 to the symmetrized bidisc and the tetrablock to recover known results from the literature (see [7, 13, 14, 20, 26]). However, as far as we know, the following formula for the Szeg kernel of the tetrablock appears to be unnoticed.
where and denotes the standard inner-product in Recall that the tetrablock first appeared in [1, DefinitionΒ 1.1], is given by
In the second half of SectionΒ 3, we apply TheoremΒ 1.3 to construct a Hardy space over the -dimensional fat Hartogs triangle
The space is the RKHS associated with the kernel given by
where we used the convention that a product over an empty set is By Aglerβs hereditary functional calculus, defines a bounded linear operator for any commuting -tuple of bounded linear operators on a Hilbert space with its Taylor spectrum contained in (refer to [2, SectionΒ 4.2]). This together with [9, TheoremΒ 2.1] leads to the following variant of von Neumannβs inequality.
Corollary 1.5.
Let be a commuting -tuple of operators in If the Taylor spectrum of is contained in the -dimensional fat Hartogs triangle and then for every
2. Proof of the Main Theorem
The proof of TheoremΒ 1.3 involves several lemmas. We begin with a variant of the removable singularity theorem for functional Hilbert spaces.
Lemma 2.1.
For a domain in let be a functional Hilbert space. Let and be nonzero. Assume that there exists a sequence in such that
| converges to in | (14) |
Then extends holomorphically to
Proof.
Since the reproducing kernel is jointly continuous111Indeed, since is separately holomorphic (by assumption), by Hartogs separate analyticity theorem, is holomorphic., the assumption (14) combined with the reproducing kernel property of (see (3)) implies that
| converges compactly to | (15) |
Let on In view of the Hartogs separate analyticity theorem (see [25]), it suffices to check that extends as a separately analytic function on To see the latter, first assume that Write where and If is nowhere vanishing in a neighborhood, say of then is holomorphic in If, for some nonnegative integer then by (15) and Weierstrass convergence theorem (see [25, TheoremΒ 1.4.20]) It now follows from the one-variable Riemann removable singularity theorem that extends holomorphically in a neighborhood of Arguing as above (in each variable), one may now show that extends as a separately analytic function. This completes the proof in case It is clear from the discussion above that the same proof works in case β
We formally introduce the following notion (weaker than -properness) as it appears naturally in the proof of TheoremΒ 1.3.
Definition 2.2.
Let be domains in A holomorphic surjection is said to be admissible if
| is singleton for every | (16) |
Remark 2.3.
Any -proper holomorphic map is admissible. This is a consequence of [23, PropositionΒ 15.1.5 & TheoremΒ 15.1.9]. There do exist admissible proper holomorphic maps of higher multiplicities. For example, for a positive integer consider the map given by
Clearly, is an admissible map. By [24, TheoremΒ 5.1], is -proper.
Remark 2.4.
Lemma 2.5.
Let be an admissible proper map and let be given. Define by if If is well-defined, then is holomorphic.
Proof.
Assume that is well-defined. Since is a proper holomorphic map, by [23, RemarkΒ 15.1.10], is holomorphic on To see that is holomorphic on by RRST (see [25, TheoremΒ 4.2.1]), it suffices to show that is continuous on . Clearly, is continuous on Now for let be a sequence in such that converges to By (16), there exists a unique such that For each integer let
Claim. converges to
Since is proper, is compact. Thus every subsequence of has a convergent subsequence. Let be subsequence of converging to some By the continuity of converges to As for every we obtain which implies that (since, by (16), has only one pre-image ). Since any bounded sequence, which is not convergent, has at least two limit points (by Bolzano-Weierstrass theorem), the claim stands verified. Since is a continuous map, by the claim above, converges to β
For a special case of the following fact, see [20, p.Β 2367].
Lemma 2.6.
Proof.
By RemarkΒ 1.4(c), defines a norm. To see that is a Hilbert space, let be a Cauchy sequence in It follows that is a Cauchy sequence in Since is a Hilbert space, there exists such that converges to in Since is nonzero (see RemarkΒ 1.4(a)), by LemmaΒ 2.1, extends as a holomorphic function on
To complete the proof, it now suffices to show that for some To see this, define by if We must check that is well-defined. For this, note the following:
if then
which is independent of the preimage
if then by the assumption (16), is singleton.
Thus, in both the cases, is well-defined. This combined with LemmaΒ 2.5 shows that with which completes the proof. β
The proof of TheoremΒ 1.3 relies on the following variant of the Waring-Lagrange theorem (cf. [2, TheoremΒ 7.1]).
Theorem 2.7.
Let be an affine-linear involution such that and let be a -invariant -proper holomorphic map. Then
Proof.
Let Since is -invariant, we have
To see the reverse inclusion, let such that Define a function by setting where and
Claim. is well-defined.
By RemarkΒ 2.3, is admissible (see DefinitionΒ 2.2). Consequently, by (16), is well-defined on If then by RemarkΒ 2.4 and the assumption that that is, or . To complete the verification of the claim, let and note that Since is -proper, Now the -invariance of yields the well-definedness of on One may now apply LemmaΒ 2.5 to complete the proof. β
Proof of TheoremΒ 1.3.
(i) To see this, let Since maps into (see RemarkΒ 1.4(d)), it suffices to check that there exists such that We claim that extends as a holomorphic function to Since by RemarkΒ 2.4, On the other hand, Thus, for any This yields the inclusion
By assumption, is affine-linear. This combined with (17) and Hartogs separate analyticity theorem shows that extends as a holomorphic function, say (for a similar reasoning, see the discussion following (15)). As and are same and -invariant outside the nowhere dense set is also -invariant. This combined with the assumptions that is a -proper map and shows that there exists such that (see TheoremΒ 2.7). It follows that
(ii) We now show that is a reproducing kernel Hilbert space (cf. [14, Lemma 3.3])). By (i), note that for any there exists such that Thus, for each extends to the holomorphic function on Further, note that for each the function belongs to Define
It is easy to see using the reproducing property of (see (3)) that
To complete the proof, it suffices to show that is the reproducing kernel for To see that, let and Since the zero set of is nowhere dense, by a routine density argument, we may assume that . Since is unitary from onto (by (i)),
This completes the proof. β
Remark 2.8.
Let be as in the statement of TheoremΒ 1.3. Assume that are multipliers of Let Then and hence by our assumption It follows that
Thus, for and for each
where denotes the operator of multiplication by This combined with TheoremΒ 1.3 yields the following:
Corollary 2.9.
Let be as in the statement of TheoremΒ 1.3. Assume that are bounded. If are multipliers of then are multipliers of In this case, are multipliers of and the commuting -tuple on is unitarily equivalent to the commuting -tuple on
3. Illustrations
In this section, we first apply TheoremΒ 1.3 to obtain Hardy spaces of the symmetrized bidisc and tetrablock. These constructions recover [20, Theorem 3.1] (in two variables) and [14, Example 3.11]. We also apply TheoremΒ 1.3 to construct Hardy spaces of the -dimensional fat Hartogs triangle and an egg domain.
3.1. Hardy space of the symmetrized bidisc
Consider the symmetrized bidisc given by
This domain first appeared in [3]. Let and Let be the involution given by Clearly, is -invariant and Consider the map given by
By [24, TheoremΒ 5.1], is -proper. Clearly, is -invariant and it maps onto . Let be the Hardy space of the unit bidisc that is, the reproducing kernel Hilbert space with the reproducing kernel
By TheoremΒ 1.3, we obtain the reproducing kernel Hilbert space Moreover, maps isometrically onto
On the other hand, for
Since it now follows from (10) that the reproducing kernel of is given by
This recovers [20, Theorem 3.1] in two variables. It can be seen that (see [5, p.Β 513], [20, CorollaryΒ 2.4]).
3.2. Hardy space of the tetrablock
Following [17, p.Β 5], the Cartan domain of type II in is given by
By [17, EqsΒ (4.5.1)&(4.7.6)], the kernel
is positive semi-definite. The Hardy space of the Cartan domain of type II is the reproducing kernel Hilbert space associated with (the equivalence of this definition with the one given on [16, p.Β 521] follows from [17, TheoremΒ 4.6.1]).
Let and (see (12)). Let be the involution given by
Since is -invariant. Clearly, Consider the map given by
By [1, Theorem 2.2], maps onto Moreover, is a -proper map (see [24, TheoremΒ 5.1]) and -invariant. Applying TheoremΒ 1.3 to with above choices of and we obtain the reproducing kernel Hilbert space Moreover, maps isometrically onto
In the following discussion, we describe the reproducing kernel of A routine calculation using
shows that
and hence, we have
Since it now follows from (10) that
To find a neat expression for consider the open set
and note that (for example, consider or ). We claim that can be expressed solely in terms over To see this, let
Note that for
This is a function in and which indeed are functions in Hence, after polarizing the real analytic function of we obtain a function holomorphic in and anti-holomorphic in This yields the formula (11) for (modulo constant).
3.3. Hardy space for the -dimensional fat Hartogs triangle
For a positive integer Let be the -dimensional Hartogs triangle in Recall from [10, 21] that the Hardy space of is the reproducing kernel Hilbert space associated with the reproducing kernel given by
Let and (see (13)). Consider the involution given by
Note that Consider the map given by
By [24, TheoremΒ 5.1], is -proper. Clearly, maps onto and it is -invariant. Applying TheoremΒ 1.3 to with above choices of and we obtain the reproducing kernel Hilbert space Moreover, maps isometrically onto Let us compute the reproducing kernel of Set Since
Since by (10), for
It is now easy to see that
We refer to as the Hardy space of the -dimensional fat Hartogs triangle This terminology may be justified in view of the following proposition.
Proposition 3.1.
For is given by
Proof.
The space has been constructed in [12, TheoremΒ 6.2] (the case in which and ).
3.4. Hardy space of an egg domain
For an integer let denote the open unit ball in Recall that the Hardy space of is the reproducing kernel Hilbert space associated with the Szeg kernel We are interested in the egg domain
(the Bergman kernel of this domain appears in [8, SectionΒ 3.2]; refer to [11] for operator theory on general egg domains).
Let and Let be as defined in (18). Consider the -proper map as given in (19). Clearly, maps onto and it is -invariant. Applying TheoremΒ 1.3 to with above choices of and we obtain the reproducing kernel Hilbert space Moreover, maps isometrically onto We check that the reproducing kernel of is given by
For simplicity, we verify this in the case Note that
Since by (10), for
which implies that
It follows that
We refer to as the Hardy space of the egg domain This may be justified in view of the following proposition.
Proposition 3.2.
For the norm of is given by
where denotes the normalized Lebesgue measure on
Proof.
Note that
Now after change of variables and we get This gives the required formula. β
As a consequence of the proposition above, one can see that the multiplier algebra of is with equality of norms. Consequently, one may apply [9, TheoremΒ 2.1] to obtain a variant of von Neumannβs inequality for
3.5. Concluding remarks
The main result of this paper (see TheoremΒ 1.3) provides a general scheme for constructing a functional Hilbert space from a known one provided the space and domain in question have symmetries in appropriate sense. Although computing kernels involved could be computationally challenging (cf. [8, 12, 26]), this transference principle is theoretically applicable in many situations. For instance, one can apply this to the Bergman space of any of the domains considered in SectionΒ 3. We leave the details to the reader. It is worth noting that TheoremΒ 1.3 is applicable in case and are unbounded domains. Recall that the Segal-Bargmann space is the Hilbert space of holomorphic functions in complex variables such that
It is a reproducing kernel Hilbert space with the kernel (see [22, SubsectionΒ 4.4.5]). Let Applying TheoremΒ 1.3 to with choices of and as defined in (18) and (19), respectively, we obtain the reproducing kernel Hilbert space As seen in earlier examples, one may check that the reproducing kernel of is given by
Acknowledgment. We convey our sincere thanks to Gargi Ghosh for a helpful discussion on the subject matter of this paper.